### v1

# LiDAR 2.0: Hierarchical Curvy Waveguide Detailed Routing for Large-Scale Photonic Integrated Circuits

Abstract
Driven by innovations in photonic computing and interconnects, photonic integrated circuit (PIC) designs advance and grow in complexity. Traditional manual physical design processes have become increasingly cumbersome. Available PIC layout tools are mostly schematic-driven, which has not alleviated the burden of manual waveguide planning and layout drawing. Previous research in PIC automated routing are largely adapted from electronic design, focusing on high-level planning and overlooking photonic-specific constraints such as curvy waveguides, bending, and port alignment. As a result, they fail to scale and cannot generate DRV-free layouts, highlighting the need for dedicated electronic-photonic design automation tools to streamline PIC physical design. In this work, we present LiDAR, the first automated PIC detailed router for large-scale designs. It features a grid-based, curvy-aware A∗ engine with adaptive crossing insertion, congestion-aware net ordering, and insertion-loss optimization. To enable routing in more compact and complex designs, we further extend our router to hierarchical routing as LiDAR 2.0. It introduces redundant-bend elimination, crossing space preservation, and routing order refinement for improved conflict resilience. We also develop and open-source a YAML-based PIC intermediate representation and diverse benchmarks, including TeMPO, GWOR, and Bennes, which feature hierarchical structures and high crossing densities. Evaluations across various benchmarks show that LiDAR 2.0 consistently produces DRV-free layouts, achieving up to 16% lower insertion loss and 7.69× speedup over prior methods on spacious cases, and 9% lower insertion loss with 6.95× speedup over LiDAR 1.0 on compact cases. Our codes are open-sourced at link.
I Introduction
In recent years, photonic integrated circuits have attracted considerable attention due to their advantages in high-speed data transmission and low power consumption. Notable advances include photonic tensor cores (PTCs) [1, 2, 3] developed for optical neural networks (ONNs) [3], as well as photonic network-on-chips (NoCs) [4, 5, 6, 7, 8] designed for high-bandwidth on-chip communication. Driven by these emerging applications, PIC designs are rapidly increasing in complexity. As shown in Fig. 1, the number of photonic components per chip is approaching the scale of 1000 and is expected to double approximately every 2.5 years [9]. This exponential growth underscores the urgent need for advanced electronic-photonic design automation (EPDA) toolchains [10, 11, 12] that can streamline layout generation, enhance productivity, and ensure design quality.
Traditionally, the physical design of PICs follows a schematic-driven approach [13], where components are placed and connected based on the circuit topology and signal path defined in the schematic. This method seeks to minimize waveguide crossings, detours, and bends to reduce insertion loss and enhance signal integrity. In certain structured designs, routing can even be manually managed, especially when the circuit features a well-organized, no-crossing topology, as seen in crossbar arrays [14], triangular or rectangular meshes [1], or binary tree structures [15]. When such designs are optimally placed with ample spacing and well-aligned ports, devices can be directly abutted or connected via simple straight waveguides, resembling the standard cell-based layout methodology commonly used in SRAM array design.
However, as PIC designs grow in complexity, significant routing challenges emerge, making automated routing increasingly essential. These challenges typically arise when: ➊ the circuit scale surpasses the limits of manual design, involving hundreds or thousands of components and nets; ➋ the circuit features a complex topology or suboptimal placement, resulting in port misalignments, excessive crossings, and severe routing congestion; ➌ the design must be adaptable across different fabrication technologies or device variants, each introducing variations in component size and behavior that impact waveguide routing; ➍ frequent design iterations are required, making manual updates time-consuming and error-prone, especially as schematic designers often lack full visibility into waveguide routing constraints or the necessity for inserting crossings. Such scenarios frequently lead to repeated back-and-forth between schematic and layout teams, particularly during layout exploration phases involving iterative placement and routing refinement.
Most existing studies primarily focus on global routing planning for PICs. A number of optical routing algorithms have been proposed for on-chip 3D system-on-package designs [16, 17, 18], with the primary goal of optimizing signal loss and overall power consumption. Tools such as PROTON [19] and PLATON [20] offer automatic placement and routing, employing a modified Lee’s algorithm for waveguide routing. In [21], further improvements are achieved by optimizing device orientation, such as flipping and rotation, to reduce insertion loss through fewer crossings. While these global routing techniques are effective in identifying low-loss paths, they often focus solely on logical path planning and tend to neglect physical realizability. As a result, they may suffer from routing congestion, infeasible bend or crossing insertions, ultimately leading to an invalid routing solution.
Some research efforts have also targeted the detailed routing stage in PIC layout design. The work [22] proposed a method using mixed integer programming for global routing, followed by Manhattan grid-based detailed routing with crossings as constraints. However, this grid-based method only supports 90∘ bends. A follow-up approach [23] introduced non-Manhattan channel routing to better accommodate curved waveguides. Despite this, modern PICs typically rely on a single optical waveguide layer, making crossings unavoidable and leaving limited flexibility for crossing optimization. To address these challenges, a fully automated PIC router is needed that is physically aware of waveguide and component instantiations and capable of performing design-rule-compliant routing with intelligent crossing insertion.
In this work, we propose LiDAR 2.0, a hierarchical PIC detailed routing tool that adopts a bottom-up routing strategy, supports non-Manhattan curvy waveguides, and enables adaptive crossing insertion. It addresses key limitations of existing methods by jointly optimizing insertion loss and enforcing layout constraints, with full awareness of waveguide, bend, and crossing geometries. Unlike schematic-driven approaches that require manual crossing insertion, LiDAR 2.0 adaptively inserts crossings during routing and produces nearly design-rule-violation (DRV)-free layouts within minutes, significantly reducing the need for post-routing fixes or iterative schematic updates. The main contributions are summarized as follows.
1.
We propose a fully automated PIC detailed routing tool, LiDAR 2.0, that generates real GDSII layouts with low insertion loss, supporting curvy waveguide geometries and automatic crossing insertion for large-scale photonic circuits within minutes.
2.
PIC Routing Benchmark: We introduce a hierarchical PIC intermediate representation (PIC IR) and open-source scalable benchmark generators for PICs, spanning diverse routing complexities, enabling realistic and challenging evaluation of PIC routing algorithms.
3.
Curvy-Aware Non-Manhattan A∗ Routing: We develop a customized A∗ search algorithm with adaptive neighbors to efficiently handle curvy structures and non-Manhattan routing patterns.
4.
Conflict-Resilient Routing: We boost routability through accessibility-enhanced port assignment, congestion and crossing space penalties, and group-based net ordering with congestion-penalized rip-up and reroute, effectively reducing access conflicts and unnecessary crossings.
5.
Extensive evaluations show that LiDAR 2.0 supports PIC routing with various crossing sizes and bend radii, achieving DRV-free layouts with up to 16% lower insertion loss (IL) and 7.69 speedup over prior methods on spacious benchmarks, and nearly DRV-free layouts with 9% lower IL and 6.95 speedup over LiDAR 1.0 on compact benchmarks.
II Preliminaries
This section first reviews related VLSI routing methods and outlines key differences from PIC routing by examining PIC-specific design rules. We then describe the conventional manual PIC routing flow, followed by routing evaluation metrics and associated challenges. Notations used in this paper are summarized in Table I.
| Symbol | Description |
| The set of nets specified in the circuit netlist. | |
| The net in , . | |
| The set of all paths. | |
| The path in , . | |
| The insertion loss of the . | |
| The maximum insertion loss over all paths. | |
| The propagation loss of the path. | |
| The crossing loss of the path. | |
| The bending loss of the path. | |
| , , | Coefficient of , , and . |
| The port group. | |
| Check region of group-based congestion penalty. | |
| Check region of crossing space penalty. | |
| The coefficient of congestion and crossing space penalty. | |
| Routing grid size. |
II-A VLSI Detailed Routing
VLSI detailed routing must address challenges such as complex design rules, pin accessibility, and limited routing resources [24]. Common routing approaches [25, 26, 27, 28] employ path-finding algorithms like A∗ search or maze routing, typically supported by a design rule checking (DRC) engine. A widely adopted strategy to resolve routing conflicts is negotiation-based routing [29], which iteratively applies rip-up and reroute techniques to eliminate routing failures.
One of the key differences between VLSI and PIC routing lies in the routing geometry. While VLSI routing typically follows Manhattan or unidirectional styles, PIC routing requires curvilinear waveguides. To support more flexible layouts, octagonal routing and other non-Manhattan styles have been explored in analog [30, 31], PCB [32, 33, 34], and package routing [35, 36]. However, research specifically targeting curvy path routing remains limited.
II-B Photonic Design Rules
PIC routing typically operates on a single waveguide layer, where all nets are 2-pin point-to-point optical paths. Below, we summarize key design rules and highlight photonic-specific considerations.
II-B1 Waveguide Spacing
Waveguides must maintain sufficient spacing from each other and from photonic device structures to prevent unwanted crosstalk. As shown in Fig. 2(a), the minimum required spacing depends on various factors such as wavelength, polarization mode, refractive index contrast, substrate type, and waveguide cross-section. For high-index contrast systems (e.g., silicon-on-insulator), relatively small spacings of 1–3 are typically sufficient.
II-B2 Bend Radius
In contrast to the sharp 90∘ bends used in VLSI metal routing, waveguides in PICs require smooth bends to minimize mode mismatch and radiation losses. To mitigate these losses, bends are typically implemented as circular arcs, Euler curves, or sine curves, depending on the routing scenario, as illustrated in Fig. 2(b).
The bend radius can vary significantly from a few microns to millimeters, depending on material properties and waveguide design. For example, silicon waveguides with high refractive index contrast support small bend radii of 5–10 , while silicon nitride waveguides, which have lower index contrast, typically require larger bend radii of 20–100 . Although larger bend radii reduce insertion loss, they also increase area consumption and may limit routing flexibility.
II-B3 Waveguide Crossing
Unlike VLSI routing, which forbids wire crossings and uses vias for layer transitions, PICs permit waveguide crossings (CRs) on the same layer, which are often necessary for high-density designs. Each crossing introduces insertion loss, typically 0.1–1 dB, and occupies an area of approximately 55 .
The crossing angle is also critical: CRs should ideally intersect at 90∘ to minimize crosstalk. This constraint complicates routing in dense layouts, as parallel waveguides require extra space to adjust their direction via bending before forming perpendicular intersections, as shown in Fig. 2(c).
II-B4 Port Connection and Alignment
Waveguides are connected via precise port abutment, which requires exact face-to-face alignment. When there is an offset between ports, additional bending is required to compensate for the misalignment, as illustrated in Fig. 2(d). Misalignment or offset at the ports can break the optical path, making precise alignment essential for successful routing. An example of correct alignment is shown in Fig. 2(e).
II-B5 Signal Integrity
Insertion loss is a critical metric for evaluating PIC routing quality, as it directly affects the laser power budget and signal integrity, including signal-to-noise ratio and crosstalk as illustrated in Fig. 2(f) The primary evaluation metric is the maximum insertion loss along the critical path. Long waveguides and excessive crossings degrade signal integrity and should be minimized wherever possible.
II-C Schematic-Driven PIC Layout
Traditional PIC layout workflows, including manual design and current EPDA tools, are schematic-driven [37]. In this approach, all structures, including crossings and waveguide segments are instantiated explicitly. Designers must plan routing during the schematic stage and manually insert crossings, while waveguide connections rely on port abutment without explicit physical net definitions.
A major limitation is that waveguide paths and crossings must be predetermined by design experts during schematic creation, based largely on empirical estimates of the final layout. Once these paths are defined, it becomes difficult to modify them during physical design, leading to a rigid and manually constrained routing topology. This inflexibility often results in repetitive revisions between the schematic and layout stages, which can be time-consuming and error-prone, and becomes impractical for large-scale designs.
To overcome these limitations, it is essential to adopt a new formulation of instances and nets that decouples schematic design from physical implementation. This will enable automated crossing insertion and routing flexibility, improving efficiency and scalability in PIC layout design.
II-D PIC Routing Quality Metrics
In addition to standard routing metrics such as wirelength, design rule violations, and runtime, a key photonic-specific metric is the insertion loss (IL), which directly affects the link power budget and signal-to-noise ratio. IL is calculated based on the optical path, which represents the trajectory of light propagation through all cascaded components from the laser source to the photodetector.
Assume a path consists of alternating instances and nets . Some instances and nets may be shared across different paths. For simplicity and due to the lack of accurate port-specific IL data in available open-source PDKs, we assume uniform IL between any input-output port pair in multi-port photonic devices. Nevertheless, port-specific ILs can be incorporated into the same formulation when available. The total insertion loss of path is defined as the sum of insertion losses from all devices and all routed waveguide segments along the path. The loss is measured in decibels (dB), following standard convention. For net-level IL, we consider three types of losses: crossing loss (), bending loss (), and propagation loss (). Therefore, we have:
where is the total waveguide length, is the number of waveguide crossings, and is the cumulative bend angle along the path. The coefficients , , and represent the insertion loss per unit waveguide length, per crossing, and per unit bend angle, respectively, and are determined by the specific photonic technology and component structures. Minimizing insertion loss is crucial to achieving the desired optical performance and signal integrity for switching, modulation, or multiplexing applications.
The maximum insertion loss across all paths, denoted as , determines the worst-case optical budget and sets the minimum required laser power to ensure reliable detection at the outputs. Therefore, serves as the primary evaluation metric for PIC routing. The routing objective is expressed as:
II-E Problem Formulation
We formulate the PIC detailed routing problem as follows:
Problem 1 (PIC Detailed Routing).
Given a set of nets and a set of placed devices , generate a routing solution for each net such that all connections are completed without design rule violations, while minimizing .
III LiDAR: Automated PIC Detailed Routing
With an input circuit netlist, as shown in Fig. 3, we present LiDAR, a detailed PIC routing framework based on a customized grid-based A∗ search. It efficiently generates curvy waveguide paths with automatic crossings insertion, aiming to minimize the maximum insertion loss, while ensuring design rules compliance. The core routing procedure consists of the following three main phases: ➊ Port Access Assignment: assigns access ports based on port orientation and local density to facilitate smooth routing and reduce congestion; ➋ Iterative Curvy-Aware Waveguide Routing: route all nets using a curvy-aware A∗ search algorithm, guided by group-based net ordering. ➌ Crossing Optimization and Route Refinement: a local rip-up and reroute mechanism is applied to resolve conflicts and optimize crossings, followed by refinement of the routing solution to eliminate redundant bendings and produce a design-rule-clean GDS layout.
III-A Accessibility-Enhanced Port Assignment
Port accessibility is one of the most critical and challenging subroutines in PIC detailed routing. Unlike unidirectional metal pins in VLSI, PICs use directional waveguide ports that require exact face-to-face orientation and precise alignment, as shown in Fig. 4(a). Improperly oriented waveguides cannot legally connect to the target port, especially when insufficient space is available to reorient the waveguide using bends. This problem becomes even more pronounced when nearby waveguides obstruct the port region. This difficulty stems from the large area needed for curvy bends. To address this, we propose port access assignment techniques that consider both port orientation and spatial density to enhance accessibility.
Port Propagation. Some waveguide ports are embedded inside the bounding box of a device. Since devices are treated as routing blockages, we propagate these internal ports outward to the device boundary based on their orientations, as illustrated in Fig. 4(a). This ensures they are reachable during routing.
Bending-Aware Port Access Region Reservation. To prevent other waveguides from blocking access to ports, a region in front of each port is reserved along its orientation as shown in Fig. 4(a). The reserved size is adaptively determined based on the waveguide’s bending radius, ensuring sufficient space for bends and improving the success rate of port access while minimizing area overhead.
Congested Port Spreading. Certain PIC devices feature dense clusters of ports that may overlap on the same routing grid, leading to access conflicts. To alleviate this, we apply a symmetric spreading strategy, where high-density ports are redistributed with a predefined spacing and extension length, as shown in Fig. 4(b). These new access ports are connected to their original locations via sine bends, ensuring they occupy distinct routing tracks and reducing congestion. The reserved access regions are also updated accordingly.
Channel Planning via Staggered Access Point Offsets. To further improve access in dense layouts, we introduce staggered access point regions, particularly beneficial for devices like multimode interference (MMI) structures that have many ports aligned on one side, as shown in Fig. 4(c). High port density may cause access ports to be blocked by nearby waveguides, while closely spaced parallel waveguides can prevent other nets from inserting crossings over them. We address this by progressively extending the access region length for inner ports and introducing an access offset larger than the crossing footprint. This approach provides sufficient space for bends and enables consecutive waveguide crossings, allowing other nets to traverse through parallel routing paths.
III-B Port-Group-based Net Order
LiDAR is a sequential router that processes one net at a time, where the routing order has a significant impact on both quality and feasibility. To address this, we propose a port-group-based net ordering strategy that clusters ports on the same device by direction, denoted as , e.g., 0∘ and 180∘ ports form groups and in Fig. 4(c) The routing proceeds group by group based on an inter-group routing order. Within each group, nets are routed in a defined intra-group order before moving to the next group. This strategy is motivated by the observation that most congestion and routing conflicts occur between nets in the same group. By routing nets group-wise with mutual awareness, intra-group conflicts are effectively reduced, improving overall routing quality.
The inter-group routing order of is first determined by its group priority score , where is the Euclidean distance of net . A smaller indicates a higher routing priority for that group. If , the group that enters the priority queue earlier will be routed first. After that, we define the net routing order within the same group.
Since ports within a group form a staggered access point region, they are more accessible and are designated as target ports, while the others serve as sources to reduce conflicts. However, the access order (i.e., intra-group routing order) is critical; an improper access sequence may lead to ports being blocked by waveguides or other port regions. Therefore, we determine the access sequence based on the source ports spatial distribution. Details are provided in Section IV.
III-C Non-Manhattan Waveguide Routing with Curvy-Aware A∗ Search
Unlike Manhattan-style VLSI routing, PICs require non-Manhattan paths to support smooth curves that reduce bending and insertion loss. We propose an iterative waveguide routing algorithm based on A∗ search, supporting both 45∘ and 90∘ turns with adaptive crossing insertion. Built on a flexible VLSI-proven framework, the algorithm enables efficient multi-objective optimization for PICs.
III-C1 Spacing-Ensured A∗ Routing Grid Size Setting
In LiDAR, the routing grid size is set to be larger than the waveguide width. Since waveguides in PICs are generally wider than their ports, this setting improves routing efficiency and simplifies port access by reducing local congestion near narrow access points.
III-C2 Parametric Curvy-Aware Neighbor Candidate Generation
To enable curvy-aware A∗ search, we propose a parametric neighbor generation scheme that leverages bending geometry and performs comprehensive design rule checks to filter out illegal candidates.
While traditional VLSI and PCB routing often use sharp 90∘ or fixed 45∘ turns, PIC routing requires smooth curves. We define each routing node by its spatial location and orientation, represented as a directional node . This directional representation is essential for maintaining alignment and orientation during port access. As illustrated in Fig. 5, neighbor candidates are determined by the node’s current orientation and a user-defined bending radius . Based on orientation, nodes are categorized into two states: Manhattan State (MS) and Non-Manhattan State (NMS). MS nodes, which align with the x- or y-axis, have five possible neighbors: one directly forward and four at angular offsets of 45∘ and 90∘. In contrast, NMS nodes, which follow diagonal trajectories, have three potential neighbors.
The exact positions of neighbor candidates are adaptively computed using the bend radius and grid size . Larger bending radii and smaller grids lead to larger step sizes. For example, for an MS node oriented at 0∘, its directly adjacent neighbor is 1 grid unit away, while the steps to reach the 90∘ and 45∘ neighbors are calculated as:
The ceiling function guarantees that there is enough spacing to accommodate bending structures. Note that unlike 45∘ diagonal neighbors in traditional 8-way A∗ search, where is equal to , our neighbors are the endpoint of the 45-degree curves, where and are not equal.
III-C3 Geometry-Aware Neighbor Legality Check
To ensure only feasible neighbor candidates are explored during A∗ search, each candidate must pass a geometry-aware legality check before being added to the priority queue. A neighbor is considered legal only if the corresponding waveguide segment does not violate any design rules upon geometric instantiation.
Hit No Obstacle: Geometry-Aware Spacing Check. If the neighbor does not intersect any obstacle, we instantiate the actual geometry of the connecting waveguide segment and perform a spacing check to verify that the path complies with design rules and is free of violations (DRVs).
Hit Routed Nets: Predictive Crossing Insertion. If the neighbor candidate intersects with an already routed waveguide (marked as an obstacle), we evaluate whether a waveguide crossing can be inserted to legally pass through it.
As shown in Fig. 6, several critical conditions must be satisfied to allow crossing insertion: ➊ Sufficient straight waveguide length: Crossings occupy a defined footprint and require perpendicular waveguide orientations. We ensure there is enough straight length and validate the orientation by checking the routing grid state at the potential crossing location. ➋ No overlap with blockages: The bounding box of the crossing must not intersect any existing obstacles to comply with the design rules. ➌ Port matching: For a valid crossing, the intersecting waveguides must align with the four ports of the crossing structure. This includes matching properties such as waveguide width, cross-section, and layer. By predictively checking all those legality conditions, our method adaptively enables crossing insertion when needed. This significantly reduces unnecessary detours and eliminates the rigidity of manually defined crossings in schematic-driven flows.
III-C4 Insertion Loss-Aware A∗ Search Cost
LiDAR adopts a customized A∗ search cost function that incorporates insertion loss considerations to improve both routing quality and algorithm efficiency. The total cost of a node in A∗ search is given by the standard formulation: , where denotes the accumulated cost from the source node to the current node , and is the heuristic estimate of the cost from node to the target node . The cost consists of two components: the insertion loss along the path to the current node, denoted as and calculated based on Eq. (1), and a group-based congestion penalty (GCP), denoted as to save routing resources for a group of nets:
where is a penalty coefficient that discourages routing too close to blockages or previously routed waveguides, preserving routing resources for unrouted nets within the same routing group, and is the number of grids that occupied by others in the check region as shown in Fig. 7. The check region is determined by the number of unrouted nets in its port group. As more nets are routed, decreases to avoid consuming extra space.
Since our method allows non-Manhattan routing, the shortest path is often diagonal. To ensure the A∗ heuristic remains valid and does not overestimate the cost, we define a customized heuristic as shown in Eq. (5):
where and represent the minimum and maximum differences in the x and y coordinates between the current node and the target node , respectively. Additionally, a penalty term proportional to the 45∘ bending loss, , is added when both and are non-zero, as this scenario requires at least one 45∘ bend. This penalty encourages paths that end in orientations better suited for port connection.
III-C5 Waveguide Instantiation
A key distinction of LiDAR compared to prior global routing methods is its geometry awareness. After determining a routing path, we instantiate the actual curvy waveguide geometry using the extrude function from GDSFactory [38], and record it on an overlapped, orientation-aware routing grid map, as illustrated in Fig. 8. This enables the A∗ search engine to treat existing waveguides as obstacles and efficiently perform waveguide spacing checks and determine valid locations for crossing insertion.
III-C6 Violated Net Removal
When the router fails to access the target port with the correct orientation, LiDAR applies a rip-up-and-reroute (RR) strategy. DRC is temporarily relaxed to allow routing progress, and any nets that violate design rules or conflict with existing paths are recorded. These violating nets are then removed and rerouted in subsequent iterations. To prevent repeated routing failures and reduce congestion, a history cost [39] is maintained and updated in a global history map before each net is ripped up. This history-based negotiation mechanism discourages routing over previously congested regions and helps distribute paths more evenly across the layout. In practice, this approach effectively resolves routing failures and improves overall success by balancing routing demands across all nets.
III-D Crossing-Waveguide Optimization
Waveguide crossings are sometimes inserted to bypass congested regions, circumvent obstructions, or avoid excessive detours. However, they not only introduce insertion loss, phase shifts, and occupy chip area, but they also cause crosstalk and potentially impact subsequent net routing. Consequently, we propose a local rip-up-and-reroute (LRR) strategy to eliminate unnecessary crossings and optimize the overall routing.
As shown in Fig. 9, LRR is triggered after a solution is found. If the initial routing solution (RS) contains no crossings, it is accepted directly. If crossings exist, possible causes include: (1) a blockage caused by another waveguide requiring a crossing, (2) a crossing chosen to bypass congestion regions, or (3) high propagation loss for non-crossing paths. To verify these, we rerun the routing with crossings disabled (NCS). If NCS succeeds, its result is compared against RS by insertion loss, and the lower-loss path is selected. If NCS fails, it indicates the net is blocked. We then check the blocking waveguide. If the blocking waveguide has never been ripped up before, it will be ripped up, as this operation will not affect CR re-insertions in subsequent iterations. This LRR process empirically optimizes the routing by balancing long waveguides and CRs to reduce overall loss.
III-E Routed Waveguide Refinement
Since the grid-based routing may cause slight misalignment between the port center and the grid center, a small offset can occur between the routed path and the port, as illustrated in Fig. 10. To resolve this, we adjust the initial and final segments of the waveguide to precisely align with the port, without altering the bend radius along the path. If direct adjustment is not feasible, a sine bend is introduced at the port to ensure proper alignment.
IV Extended Conflict-Resilient Hierarchical PIC Routing Scheme
To enable a more compact layout and ensure successful routing under limited routing resources, it is essential not only to minimize bending and crossing but also to reserve sufficient space for them, as they are area-consuming. This becomes even more challenging in compact layouts. To address this, we further extend our framework to LiDAR 2.0, enhancing its conflict resolution capability and enabling successful routing even under tight layout constraints.
IV-A Hierarchical Netlist Tree Construction
PIC netlists often exhibit inherent hierarchical structure, where complex systems are composed of repeated functional modules as shown in Fig. 11. Leveraging this structure is crucial for scalable physical design. To exploit such hierarchy, we first construct a Hierarchical Tree based on the invocation relationships between modules and the overall circuit. This hierarchical representation enables global routing planning at multiple levels. We then perform routing in a bottom-up manner, completing routing for each level of modules progressively. This multi-level routing approach helps avoid conflicts between different modules, as routing is localized within each submodule before higher-level integration. Similar to standard-cell design in VLSI, the internal routing resources of a module are treated as private and should not be accessed by external nets.
To further improve efficiency, we reuse routing results of identical module types via translation, rotation, and flipping. Before reuse, we apply design rule checking at the target location to ensure legality. If a reused result violates constraints, we fall back to re-routing that specific instance from scratch. This reuse mechanism significantly accelerates layout generation for large-scale, hierarchical PIC designs.
IV-B Routing Conflict Reduction via Spatial-Aware Optimization
Offset Neighbor. Although the algorithm naturally discourages excessive bending through its bend cost heuristic, it may still introduce unnecessary bends when ports are slightly misaligned. Especially in tight spaces, such misalignment can trigger multiple bends, increasing insertion loss and risking routing failure, as shown in Fig. 12(a).
To address this issue, we incorporate additional offset neighbors to avoid such detours under specific conditions: when the current node is close to the target node (within four times the bending radius), and the orientation differs by 180 degrees from the target node. These offset neighbors are designed to correct small misalignments (less than the bending radius) between the source and target ports. The offset neighbor locations are computed analytically based on the predefined bending radius and grid size, ensuring that the resulting path remains smooth and physically realizable using sine-bend waveguides. From an algorithmic perspective, the introduction of offset neighbors simplifies port access, reduces redundant neighbor exploration caused by bending detours, and thereby improves the search efficiency.
Preserved Crossing Space. A key source of routing conflict in PICs is the limited space for inserting crossings. Unlike VLSI vias, which can be inserted independently of other nets, photonic waveguide crossings must be placed directly on top of the crossed waveguides and require significant area, making insertion difficult in dense regions. As shown in Fig. 12(b), closely spaced waveguides may block required consecutive crossings, causing routing failure.
To mitigate the crossing space conflict, we introduce a crossing space penalty during the A∗ search to encourage the router to preserve sufficient spacing around the nets that are likely to be crossed, as inferred from the netlist topology. The penalty is defined as:
where is a user-defined penalty coefficient and denotes the minimum width required to insert a valid crossing structure. The penalty is applied over a -wide region around the net segment, discouraging other nets from occupying this reserved space.
IV-C Reduce Routing Conflict via Routing Order Optimization
Although the port region allows access points to be staggered, reducing conflicts during access, an inappropriate access order may easily result in the target ports being blocked by routed waveguides and other port regions. Especially for multi-port components, the intra-group routing order is critically important. As shown in Fig. 13(a), both the orientation and spatial location of ports significantly influence the routing order, as they jointly determine the accessibility and potential for routing conflicts among nets.
To address the intra-group routing order, we determine the routing order by the spatial relationship between source ports and target ports within the group, as shown in Fig. 13(b). Let the set of source ports be denoted as . We compute their geometric mean coordinates. If each port has a position , the average position is:
After obtaining the mean location of the source ports, we determine the routing order based on the Manhattan distance from each target port to . Target ports that are closer to the average source location are assigned higher routing priority. This approach can effectively resolve most routing conflict issues, particularly in multiport components.
Routing Order Refinement. It is difficult to guarantee an optimal routing order. In particular, when nets are routed independently without group information, the aforementioned method becomes inapplicable. Hence, it is necessary to dynamically adjust the order during the routing process. An inappropriate order may lead to ports being blocked by previously routed waveguides, resulting in escape or access failures. As illustrated in Fig. 14, since net1, net2, and net3 are routed independently, net2 fails because it cannot access the port due to net1’s port region and net3’s waveguide blocking its path.
To address routing failures, we enhance the rip-up and re-route scheme by not only removing nets that directly conflict with the failed path but also ripping up those located in front of the failed net’s ports to alleviate congestion. As shown in Fig. 14, we scan along the port orientation and identify the conflict order and names of the conflicting nets (starting from ), which serve as guidance to adjust the routing order. Higher-priority nets are routed first to improve escape success, but may block others’ access. To avoid this, we dynamically extend the target port regions of high-priority nets, shifting their access points and reducing the risk of subsequent blocking.
Our extended conflict-resilient hierarchical routing scheme can effectively enhance routability even under tight layout constraints and help accelerate the routing process.
V PIC Routing Benchmark
| Benchmarks | # Components | # Nets | # Total CR | Die Size () |
| Clements_88 | 52 | 79 | 0 | 4800 1600 |
| Clements_1616 | 168 | 287 | 0 | 8000 3200 |
| ADEPT_88 | 82 | 111 | 33 | 4400 1600 |
| ADEPT_1616 | 162 | 223 | 55 | 6900 3200 |
| ADEPT_3232 | 318 | 446 | 121 | 13000 6400 |
| Light_(a,b,c,d) | 9 | 16 | - | 10000 10000 |
| Benchmarks | # Components | # Nets | # Total CR | Die Size () |
| Clements_88_C | 60 | 87 | 0 | 3500 1000 |
| Clements_1616_C | 184 | 303 | 0 | 6800 2000 |
| ADEPT_88_C | 104 | 127 | 30 | 4400 1000 |
| ADEPT_1616_C | 227 | 319 | 80 | 6200 2000 |
| ADEPT_3232_C | 529 | 767 | 160 | 7800 4000 |
| TeMPO_88_C | 452 | 515 | 49 | 1700 1500 |
| TeMPO_1616_C | 1796 | 2051 | 225 | 3000 3400 |
| TeMPO_3232_C | 7172 | 8195 | 961 | 5700 6500 |
| GWOR_1616_C | 17 | 32 | – | 4000 4000 |
| GWOR_3232_C | 33 | 64 | – | 6000 6000 |
| Bennes_1616_C | 224 | 416 | 88 | 2200 1300 |
| Bennes_3232_C | 576 | 1024 | 416 | 3500 2500 |
PIC Intermediate Representation. Inspired by the LEF/DEF formats widely used in electronic design automation, we introduce a YAML-based photonic intermediate representation (PIC IR) to describe the connectivity and types of photonic components. This IR supports hierarchical module definitions, enabling modular reuse and streamlined management of complex circuits. To bridge this abstract representation with physical layouts, we develop a translation tool that interfaces with GDSFactory, allowing for both visualization and generation of valid GDSII layouts from the IR.
Benchmark Suites. Built upon the PIC IR, we further develop and open-source a suite of parameterized benchmark generators. These scripts enable users to generate benchmarks with configurable scales, such as the number of components, number of nets, and overall circuit dimensions. They also support fine-grained control over routing complexity by adjusting parameters including component pitch (i.e., component density), port density, port misalignment, and the number of topological crossings, allowing for the creation of benchmarks with varying levels of routing difficulty. The benchmark suite includes both photonic tensor core and optical switch, providing a diverse set of testcases to evaluate the router in realistic PIC settings. Users can also define their own PIC benchmarks easily by using our interface.
For the photonic tensor core, we provide the following benchmarks, each available at multiple scales:
Clements [40]: A highly structured Mach-Zehnder interferometer (MZI) array with a regular topology and no topological crossings. It is relatively straightforward to route when sufficient layout space is available.
ADEPT [41]: An auto-searched photonic tensor core that includes numerous multiport components and a large number of inherent crossings, resulting in high routing congestion and increased difficulty.
TeMPO [42]: A time-multiplexed dynamic photonic tensor accelerator with a hierarchical crossbar structure. It contains multiple computing nodes and large fan-out MMI components, along with a significant number of crossings, making routing particularly challenging.
For the optical switch category, we use benchmarks below:
Light [43]: Based on macro-ring resonators (MRRs), it features an unstructured topology with a centralized MRR macro. We define four layout variants according to memory controller and hub placement to evaluate routing under different traffic patterns, focusing on minimizing crossings and insertion loss.
GWOR [44]: Also MRR-based, it splits the central switch into multiple distributed parts, with I/O ports placed uniformly along the chip periphery and no group-level planning. This benchmark emphasizes routing conflict resolution via pure spatial reasoning.
Bennes [45]: Based on Mach-Zehnder interferometers, it adopts a structured, hierarchical Benes topology. It exhibits extremely high crossing density, where a single net may require tens of crossings, challenging both crossing reservation and congestion-aware routing strategies.
Tables II and III summarize the benchmark statistics. Total #CR refers to the number of topological crossings inherent to the circuit and resolved during routing. For Light and GWOR, internal switch crossings are excluded, as we focus solely on the routing between the switch and the I/O interfaces in these benchmarks. Table II (ISPD version) includes spacious, moderately complex layouts, while Table III features more compact benchmarks with higher component density, smaller die sizes, and more crossings, posing greater routing challenges and better reflecting real-world photonic layout constraints. The bending radius is set to 60 for Light and GWOR and 5 for others.
| Benchmarks | PROTON [19] (Adaptive crossing penalty) | LiDAR 1.0 [46] | LiDAR 2.0 | ||||||||||||
| #CR | WL | #DRV | Time (s) | #CR | WL | #DRV | Time (s) | #CR | WL | #DRV | Time (s) | ||||
| Clements_88 | 0 | 3.39 | 16.99 | 0 | 112 | 0 | 2.94 | 16.38 | 0 | 29 | 0 | 2.89 | 15.98 | 0 | 7 |
| Clements_1616 | 5 | 5.06 | 29.31 | 12 | 527 | 0 | 4.38 | 26.74 | 0 | 144 | 0 | 4.07 | 26.03 | 0 | 61 |
| ADEPT_88 | 16 | 4.70 | 17.12 | 26 | 194 | 18 | 4.10 | 18.00 | 0 | 71 | 18 | 3.99 | 17.63 | 0 | 65 |
| ADEPT_1616 | 28 | 7.84 | 24.07 | 98 | 1395 | 16 | 7.38 | 17.80 | 0 | 243 | 16 | 6.95 | 17.20 | 0 | 243 |
| ADEPT_3232 | 66 | 16.13 | 44.57 | 355 | 10894 | 50 | 15.04 | 36.34 | 0 | 1348 | 50 | 14.70 | 36.06 | 0 | 1038 |
| Light_a | 12 | 32.98 | 11.09 | 0 | 39 | 0 | 31.11 | 7.78 | 0 | 101 | 0 | 29.99 | 7.61 | 0 | 103 |
| Light_b | 6 | 18.71 | 5.89 | 0 | 8 | 0 | 21.55 | 6.31 | 0 | 44 | 0 | 21.12 | 6.24 | 0 | 47 |
| Light_c | 14 | 20.81 | 10.23 | 1 | 52 | 0 | 35.29 | 8.40 | 0 | 72 | 0 | 35.92 | 8.57 | 0 | 74 |
| Light_d | 13 | 28.49 | 10.94 | 1 | 53 | 0 | 33.52 | 8.14 | 0 | 80 | 0 | 32.82 | 8.09 | 0 | 83 |
| Geo-mean | - | 15.35 | 18.91 | - | 1475 | - | 17.26 | 16.21 | - | 237 | - | 16.94 | 15.93 | - | 191 |
| Ratio | - | 1 | 1 | - | 1 | - | 1.12 | 0.86 | - | 0.16 | - | 1.1 | 0.84 | - | 0.13 |
VI Evaluation Results
The proposed photonic detailed routing framework is implemented in Python based on GDSFactory 8 [38] libraries. All evaluations are conducted on a on a Linux server with a 128-core AMD EPYC 7763 CPU and 512-GB memory. Note that runtime and results may slightly differ from the ISPD version due to platform and GDSFactory version variations.
The placement solutions of all benchmark circuits are verified with simulation using GDSFactory and KLayout. Device insertion losses used for critical path IL evaluation are listed in Table V, with bending loss proportional to angle.
Baselines. We compare LiDAR 2.0 with PROTON [19], which performs path planning with adaptive crossing penalties but lacks valid waveguide geometry generation. For fair comparison, we adapt PROTON with reserved port regions and global rip-up and reroute support. If no feasible solution is found due to congestion, DRC constraints are temporarily relaxed for specific failed nets for meaningful comparison of #DRVs. LiDAR 1.0 is also included to demonstrate the improvements in conflict handling, crossing insertion, and runtime efficiency in LiDAR 2.0.
VI-A Results on Spacious Benchmarks
We first evaluate the routing algorithm under spacious layout conditions. Table IV presents a comparison of routing metrics on the spacious benchmarks. In this setting, LiDAR 1.0 and LiDAR 2.0 achieve comparable performance and successfully generate DRV-free layouts across all benchmarks. Compared to PROTON, LiDAR 2.0 achieves an average of 16% lower critical path IL and 7.69 speedup.
Analysis of PTC Results. Despite the Clements circuit having a regular topology with no inherent crossings and relatively spacious layouts, insufficient planning and lack of reserved access space lead to unnecessary crossings and DRVs in PROTON. These routing conflicts become more severe in ADEPT benchmarks, which features dense multi-port MMIs and a large number of crossings. By incorporating Accessibility-Enhanced Port Assignment and Port-Group-Based Net Ordering, LiDAR 2.0 effectively mitigates these conflicts and reduces unnecessary rip-up and reroute. As circuit scale increases, PROTON suffers from escalating #DRVs and runtime, while LiDAR 2.0 consistently produces DRV-free, low-IL layouts.
Analysis of Optical Switch Results. Light benchmark provides abundant routing resources and is primarily used to evaluate the router’s ability to optimize insertion loss and crossings. Notably, there exists a counter-intuitive trade-off between #CR and WL, minimizing #CR may lead to long detours that increase propagation loss, resulting in a higher overall . Except for the Light_b case, LiDAR achieves the lowest across all other cases while maintaining crossing-optimal (#CR=0), DRV-free layouts.
VI-B Results on Compact Benchmarks
| Benchmarks | PROTON [19] (Adaptive crossing penalty) | LiDAR 1.0 [46] | LiDAR 2.0 | ||||||||||||
| #CR | WL | #DRV | Time (s) | #CR | WL | #DRV | Time (s) | #CR | WL | #DRV | Time (s) | ||||
| Clements_88_C | 0 | 2.12 | 16.28 | 2 | 11 | 0 | 1.91 | 16.23 | 2 | 11 | 0 | 1.91 | 16.23 | 0 | 9 |
| Clements_1616_C | 1 | 3.91 | 27.13 | 5 | 93 | 0 | 3.56 | 26.56 | 5 | 101 | 0 | 3.56 | 26.56 | 0 | 89 |
| ADEPT_88_C | 14 | 4.53 | 16.26 | 11 | 96 | 14 | 3.71 | 16.12 | 2 | 82 | 14 | 3.71 | 16.12 | 0 | 68 |
| ADEPT_1616_C | 34 | 8.24 | 27.60 | 42 | 515 | 21 | 8.63 | 21.09 | 7 | 304 | 21 | 8.35 | 21.02 | 0 | 236 |
| ADEPT_3232_C | 49 | 10.78 | 36.20 | 171 | 1496 | 31 | 10.58 | 27.07 | 30 | 735 | 30 | 9.83 | 26.48 | 3 | 568 |
| TeMPO_88_C | 12 | 1.73 | 18.48 | 14 | 95 | 10 | 1.90 | 17.33 | 5 | 147 | 7 | 2.83 | 15.86 | 0 | 44 |
| TeMPO_1616_C | 27 | 3.47 | 34.58 | 35 | 251 | 23 | 3.86 | 32.27 | 24 | 1571 | 15 | 5.48 | 28.31 | 0 | 119 |
| TeMPO_3232_C | 49 | 5.91 | 62.57 | 113 | 1125 | 46 | 6.11 | 60.45 | 57 | 8057 | 31 | 10.96 | 53.27 | 0 | 425 |
| GWOR_1616_C | 30 | 3.09 | 16.72 | 35 | 19 | 27 | 2.78 | 15.02 | 26 | 15 | 21 | 2.80 | 12.03 | 0 | 7 |
| GWOR_3232_C | 62 | 7.22 | 35.26 | 126 | 60 | 56 | 6.81 | 31.26 | 44 | 45 | 44 | 7.82 | 25.40 | 0 | 11 |
| Bennes_1616_C | 33 | 4.42 | 22.69 | 76 | 721 | 26 | 4.28 | 18.84 | 3 | 175 | 23 | 3.57 | 17.15 | 0 | 33 |
| Bennes_3232_C | 56 | 9.24 | 36.76 | 365 | 5117 | 68 | 9.08 | 42.39 | 78 | 1690 | 58 | 6.51 | 36.56 | 0 | 251 |
| Geo-mean | - | 5.39 | 29.21 | - | 800 | - | 5.27 | 27.05 | - | 1078 | - | 5.61 | 24.58 | - | 155 |
| Ratio | - | 1 | 1 | - | 1 | - | 0.97 | 0.92 | - | 1.33 | - | 1.04 | 0.84 | - | 0.19 |
We further evaluate the routing performance on compact benchmarks characterized by higher component density and more topological crossings. For these cases, only a single round of rip-up and reroute is applied, as limited layout space makes additional iterations ineffective. Figure 15 shows the final routed layouts generated by LiDAR 2.0.
As shown in Table VI, LiDAR 2.0 achieves nearly DRV-free results across all benchmarks, with an average 16% reduction in insertion loss and a 5.16 speedup over PROTON, and a 6.95 speedup and 9% lower insertion loss compared to LiDAR 1.0. On the TeMPO benchmark, PROTON shows a much lower runtime due to complete port blocking and early termination without successful routing.
On the compact cases, LiDAR 1.0 also shows increased DRVs and runtime, especially on newly added cases. In contrast, LiDAR 2.0 adopts a hierarchical routing strategy that enables subcircuit reuse in designs such as TeMPO and Bennes, significantly improving runtime. Building on this, intra-group routing order optimization and proactive crossing space preservation further reduce conflicts in multiport components, improve solution quality, and avoid unnecessary search. As a result, LiDAR 2.0 successfully achieves DRV-free routing even on high-crossing-density Bennes benchmarks. For GWOR, which lacks group information, LiDAR 2.0 still generates DRV-free results by dynamically refining the routing order and port regions during rip-up and reroute.
Figure 16 presents the runtime breakdown across various routing stages on Bennes_3232. The substantial speedup achieved by LiDAR 2.0 stems from its ability to minimize routing conflicts, eliminate redundant searches and RR iterations, and hierarchically reuse routing results.
VI-C Results on Different Crossing Sizes and Bend Radii
| Benchmarks | Crossing size = | Crossing size = | Crossing size = | ||||||||||||
| #CR | WL | #DRV | Time (s) | #CR | WL | #DRV | Time (s) | #CR | WL | #DRV | Time (s) | ||||
| ADEPT_1616_C | 21 | 8.35 | 21.02 | 0 | 236 | 21 | 8.54 | 21.08 | 0 | 299 | 21 | 9.08 | 20.96 | 3 | 379 |
| ADEPT_1616_L | 21 | 8.84 | 21.05 | 0 | 339 | 21 | 8.80 | 21.06 | 0 | 347 | 21 | 9.18 | 21.13 | 1 | 373 |
| TeMPO_1616_C | 15 | 5.48 | 28.31 | 0 | 119 | 15 | 5.41 | 28.30 | 0 | 127 | 15 | 5.33 | 28.29 | 0 | 131 |
| TeMPO_1616_L | 15 | 4.86 | 28.22 | 0 | 132 | 15 | 4.79 | 28.21 | 0 | 136 | 15 | 4.71 | 28.19 | 0 | 139 |
| Bennes_1616_C | 23 | 3.57 | 17.15 | 0 | 33 | 23 | 3.45 | 17.13 | 0 | 34 | 23 | 3.34 | 17.15 | 0 | 34 |
| Bennes_1616_L | 22 | 3.70 | 16.68 | 0 | 34 | 22 | 3.57 | 16.62 | 0 | 28 | 23 | 3.49 | 17.23 | 0 | 35 |
| Metrics | GWOR_3232 | Clements_1616_C | ADEPT_1616_C | Bennes_1616_C | TeMPO_1616_C | ||||||||
| =10 | =20 | =40 | =80 | =100 | =10 | =15 | =10 | =15 | =5 | =10 | =5 | =10 | |
| WL | 7.77 | 7.77 | 7.79 | 7.85 | 8.06 | 3.49 | 3.33 | 8.55 | 8.52 | 3.57 | 3.76 | 2.83 | 3.07 |
| #CR | 44 | 44 | 44 | 44 | 44 | 0 | 3 | 21 | 22 | 23 | 23 | 7 | 15 |
| 26.40 | 25.40 | 25.40 | 25.41 | 25.44 | 26.54 | 26.25 | 21.06 | 21.62 | 17.15 | 17.32 | 15.86 | 27.90 | |
| #DRV | 0 | 0 | 0 | 0 | 3 | 0 | 3 | 0 | 2 | 0 | 0 | 0 | 0 |
| Time (s) | 10 | 10 | 10 | 14 | 16 | 96 | 111 | 298 | 339 | 33 | 33 | 44 | 172 |
Different PIC technologies impose varying constraints on crossing sizes and bending radii. Table VII and Table VIII show the routing results of LiDAR 2.0 under different crossing sizes and bend radii. Benchmarks with an L suffix indicate larger component pitch. As expected, larger crossing sizes and bend radii increase area usage and routing difficulty. Still, LiDAR 2.0 consistently generates DRV-free layouts across compact and spacious benchmarks, with only minor DRVs in ADEPT cases. LiDAR 2.0 supports the insertion of various crossing sizes and remains robust under different bend radii. It successfully routes GWOR_3232 even with large bend radii up to 80 , which is typical in silicon nitride platforms. For silicon photonics, where the bending radius is typically 5–10 , LiDAR 2.0 performs reliably. Feasible solutions remain available at 15 despite minor DRVs, except for the TeMPO and Bennes benchmarks, where no valid routing solution exists due to insufficient spacing.
VI-D Discussion
| Metrics | High Crossing Cost | Low Crossing Cost | ||
| w/o GCP | LiDAR | w/o GCP | LiDAR | |
| #CR | 6 | 0 | 5 | 5 |
| WL (mm) | 20.72 | 29.99 | 25.11 | 26.04 |
| 15.21 | 10.61 | 7.18 | 7.31 | |
| DRV | 0 | 0 | 1 | 0 |
| Time (s) | 129 | 103 | 261 | 197 |
| Benchmarks | w/o Offset Neighbor | w/ Offset Neighbor | ||||
| #DRV | Time (s) | #DRV | Time (s) | |||
| TeMPO_88_C+ | 18.00 | 22 | 96 | 15.86 | 0 | 44 |
| TeMPO_88_C | 16.02 | 0 | 67 | 15.14 | 0 | 40 |
| TeMPO_1616_C+ | 34.12 | 71 | 470 | 28.03 | 0 | 144 |
| TeMPO_1616_C | 28.58 | 0 | 215 | 28.31 | 0 | 119 |
| TeMPO_3232_C+ | 58.57 | 157 | 1713 | 52.28 | 0 | 494 |
| TeMPO_3232_C | 53.94 | 0 | 673 | 53.27 | 0 | 425 |
Crossing-Waveguide Optimization. As shown in Table IV (PROTON [19] vs. LiDAR 2.0), our proposed crossing optimization strategy introduces an extra runtime penalty on the Light benchmark, but it leads to higher solution quality as it reduces the crossings between IOs and the switch given the high crossing penalty.
Port-Group-based Congestion Penalty (GCP). We evaluate the effectiveness of our proposed group-based congestion penalty in guiding crossing optimization under different crossing insertion loss settings, as shown in Table IX. Specifically, we compare scenarios with high crossing IL () and low crossing IL (). When the crossing IL is high, our method effectively avoids crossings to minimize the maximum insertion loss (). In contrast, under low crossing IL, it prefers shorter waveguide paths with more crossings to further optimize .
Without the group-based penalty, however, the router tends to use more crossings as increases. This is because, under congestion from previously routed nets, it struggles to find low-crossing paths. By introducing the group-based penalty, routing conflicts are significantly reduced, enabling more efficient decisions with fewer crossings and lower overall IL. As illustrated in Fig. 17, the coefficient functions as a flexible control parameter, allowing users to balance between waveguide length and the number of crossings based on specific PIC performance requirements.
Offset Neighbor. Offset neighbors help eliminate unnecessary bending detours caused by port misalignment in PICs, thereby reducing both routing conflicts and search overhead. To further demonstrate their effectiveness, we evaluate the router on a compact TeMPO_C+ variant with x/y pitch reduced from 25/50 to 20/15 , as shown in Table X. In such highly constrained layouts, detours without offset neighbors often result in design rule violations. Even in more spacious cases, enabling offset neighbors consistently leads to faster runtime compared to routing without them.
VII Conclusion
In this work, we first introduce LiDAR, an open-source automated detailed routing tool specifically designed for photonic integrated circuits. LiDAR features a non-Manhattan curvy-aware A∗ search engine with accessibility-enhanced port assignment, adaptive crossing insertion, congestion-aware group-based net ordering and objective, and crossing optimization scheme to handle unique PIC routing constraints while optimizing critical path insertion loss. To further alleviate routing conflicts and improve efficiency, we extend LiDAR into a more scalable and conflict-resilient framework, LiDAR 2.0, by introducing hierarchical routing reuse, intra-group net ordering, offset-neighbor search, and crossing-space preservation. These enhancements substantially improve routability under tight layout constraints and high crossing density. LiDAR 2.0 achieves nearly DRV-free layouts across challenging benchmarks, delivering up to 9% lower insertion loss and 6.95 speedup compared to LiDAR 1.0, advancing the automation frontier for large-scale PIC physical design.

### v2

@@
 Available PIC layout tools are mostly schematic-driven, which has not alleviated the burden of manual waveguide planning and layout drawing.
-Previous research in PIC automated routing are largely adapted from electronic design, focusing on high-level planning and overlooking photonic-specific constraints such as curvy waveguides, bending, and port alignment.
+Previous research in PIC automated routing is largely adapted from electronic design, focusing on high-level planning and overlooking photonic-specific constraints such as curvy waveguides, bending, and port alignment.
 As a result, they fail to scale and cannot generate DRV-free layouts, highlighting the need for dedicated electronic-photonic design automation tools to streamline PIC physical design.
@@
 We also develop and open-source a YAML-based PIC intermediate representation and diverse benchmarks, including TeMPO, GWOR, and Bennes, which feature hierarchical structures and high crossing densities.
-Evaluations across various benchmarks show that LiDAR 2.0 consistently produces DRV-free layouts, achieving up to 16% lower insertion loss and 7.69× speedup over prior methods on spacious cases, and 9% lower insertion loss with 6.95× speedup over LiDAR 1.0 on compact cases.
+Evaluations across various benchmarks show that LiDAR 2.0 produces nearly DRV-free layouts, achieving up to 16% lower insertion loss and 7.69× speedup over prior methods on spacious cases, and 9% lower insertion loss with 6.95× speedup over LiDAR 1.0 on compact cases.
 Our codes are open-sourced at link.
@@
 In recent years, photonic integrated circuits have attracted considerable attention due to their advantages in high-speed data transmission and low power consumption.
-Notable advances include photonic tensor cores (PTCs) [1, 2, 3] developed for optical neural networks (ONNs) [3], as well as photonic network-on-chips (NoCs) [4, 5, 6, 7, 8] designed for high-bandwidth on-chip communication.
+Notable advances include photonic tensor cores (PTCs) [1] developed for optical neural networks (ONNs) [2], as well as photonic network-on-chips (NoCs) [3] designed for high-bandwidth on-chip communication.
 Driven by these emerging applications, PIC designs are rapidly increasing in complexity.
@@
 1, the number of photonic components per chip is approaching the scale of 1000 and is expected to double approximately every 2.5 years [4].
-This exponential growth underscores the urgent need for advanced electronic-photonic design automation (EPDA) toolchains [10, 11, 12] that can streamline layout generation, enhance productivity, and ensure design quality.
+This growth highlights the urgent need for advanced electronic-photonic design automation (EPDA) toolchains [5] to streamline layout, boost productivity, and ensure quality.
 Traditionally, the physical design of PICs follows a schematic-driven approach [6], where components are placed and connected based on the circuit topology and signal path defined in the schematic.
@@
 Most existing studies primarily focus on global routing planning for PICs.
-A number of optical routing algorithms have been proposed for on-chip 3D system-on-package designs [16, 17, 18], with the primary goal of optimizing signal loss and overall power consumption.
+A number of optical routing algorithms have been proposed for on-chip 3D system-on-package designs [10, 11], with the primary goal of optimizing signal loss and overall power consumption.
 Tools such as PROTON [12] and PLATON [13] offer automatic placement and routing, employing a modified Lee’s algorithm for waveguide routing.
@@
 5.
-Extensive evaluations show that LiDAR 2.0 supports PIC routing with various crossing sizes and bend radii, achieving DRV-free layouts with up to 16% lower insertion loss (IL) and 7.69 speedup over prior methods on spacious benchmarks, and nearly DRV-free layouts with 9% lower IL and 6.95 speedup over LiDAR 1.0 on compact benchmarks.
+Evaluations show that LiDAR 2.0 supports PIC routing with various crossing sizes and bend radii, achieving DRV-free layouts with 16% lower insertion loss (IL) and 7.69 speedup over prior methods on spacious benchmarks, and nearly DRV-free layouts with 9% lower IL and 6.95 speedup over LiDAR 1.0 on compact benchmarks.
 II Preliminaries
@@
 II-B Photonic Design Rules
-PIC routing typically operates on a single waveguide layer, where all nets are 2-pin point-to-point optical paths. Below, we summarize key design rules and highlight photonic-specific considerations.
-II-B1 Waveguide Spacing
-Waveguides must maintain sufficient spacing from each other and from photonic device structures to prevent unwanted crosstalk.
-As shown in Fig.
-2(a), the minimum required spacing depends on various factors such as wavelength, polarization mode, refractive index contrast, substrate type, and waveguide cross-section.
-For high-index contrast systems (e.g., silicon-on-insulator), relatively small spacings of 1–3 are typically sufficient.
+PIC routing typically operates on a single waveguide layer, where all nets are 2-pin optical paths. We summarize key design rules and highlight photonic-specific considerations.
+II-B1 Waveguide Type
+PICs often contain multiple types of waveguides, whose characteristics depend on factors such as wavelength, polarization mode, substrate type, and cross-section, as shown in Fig.
+2(a).
+Waveguide connections must be strictly matched in type or transitioned via tapers to minimize loss.
+Different waveguide types also have different spacing requirements to prevent unwanted crosstalk.
+For high-index contrast systems (e.g., silicon-on-insulator), relatively small spacings of 1–3 m are typically sufficient.
 II-B2 Bend Radius
@@
 II-B5 Signal Integrity
-Insertion loss is a critical metric for evaluating PIC routing quality, as it directly affects the laser power budget and signal integrity, including signal-to-noise ratio and crosstalk as illustrated in Fig.
+Insertion loss is a critical metric for evaluating PIC routing quality, as it directly affects the laser power budget and signal integrity, including signal-to-noise ratio and crosstalk, as illustrated in Fig.
 2(f) The primary evaluation metric is the maximum insertion loss along the critical path.
@@
 Traditional PIC layout workflows, including manual design and current EPDA tools, are schematic-driven [30].
-In this approach, all structures, including crossings and waveguide segments are instantiated explicitly.
+In this approach, all structures, including crossings and waveguide segments, are instantiated explicitly.
 Designers must plan routing during the schematic stage and manually insert crossings, while waveguide connections rely on port abutment without explicit physical net definitions.
@@
 With an input circuit netlist, as shown in Fig.
-3, we present LiDAR, a detailed PIC routing framework based on a customized grid-based A∗ search.
-It efficiently generates curvy waveguide paths with automatic crossings insertion, aiming to minimize the maximum insertion loss, while ensuring design rules compliance.
-The core routing procedure consists of the following three main phases: ➊ Port Access Assignment: assigns access ports based on port orientation and local density to facilitate smooth routing and reduce congestion; ➋ Iterative Curvy-Aware Waveguide Routing: route all nets using a curvy-aware A∗ search algorithm, guided by group-based net ordering.
-➌ Crossing Optimization and Route Refinement: a local rip-up and reroute mechanism is applied to resolve conflicts and optimize crossings, followed by refinement of the routing solution to eliminate redundant bendings and produce a design-rule-clean GDS layout.
+3, we present LiDAR, a detailed PIC routing framework based on customized grid-based A∗ search.
+It efficiently generates curvy waveguide paths with automatic crossing insertion, minimizing maximum insertion loss while ensuring design-rule compliance.
+The core routing flow has three phases: ➊ Port Access Assignment: assigns ports by orientation and local density to ease routing and reduce congestion; ➋ Iterative Curvy-Aware Routing: routes all nets with a curvy-aware A∗ guided by group-based net ordering; ➌ Crossing Optimization and Refinement: applies local rip-up and reroute to resolve conflicts and optimize crossings, then refines results to remove redundant bends and produce a clean GDS layout.
 III-A Accessibility-Enhanced Port Assignment
@@
 Port Propagation.
-Some waveguide ports are embedded inside the bounding box of a device.
+Some waveguide ports are inside the bounding box of a device.
 Since devices are treated as routing blockages, we propagate these internal ports outward to the device boundary based on their orientations, as illustrated in Fig.
@@
 This ensures they are reachable during routing.
-Bending-Aware Port Access Region Reservation.
-To prevent other waveguides from blocking access to ports, a region in front of each port is reserved along its orientation as shown in Fig.
-4(a).
-The reserved size is adaptively determined based on the waveguide’s bending radius, ensuring sufficient space for bends and improving the success rate of port access while minimizing area overhead.
 Congested Port Spreading.
@@
 4(b).
-These new access ports are connected to their original locations via sine bends, ensuring they occupy distinct routing tracks and reducing congestion.
-The reserved access regions are also updated accordingly.
-Channel Planning via Staggered Access Point Offsets.
-To further improve access in dense layouts, we introduce staggered access point regions, particularly beneficial for devices like multimode interference (MMI) structures that have many ports aligned on one side, as shown in Fig.
-4(c).
-High port density may cause access ports to be blocked by nearby waveguides, while closely spaced parallel waveguides can prevent other nets from inserting crossings over them.
-We address this by progressively extending the access region length for inner ports and introducing an access offset larger than the crossing footprint.
-This approach provides sufficient space for bends and enables consecutive waveguide crossings, allowing other nets to traverse through parallel routing paths.
+These new access ports are connected to their original locations via sine bends.
+We add 5 units of extension length for each grid shift, ensuring they occupy distinct routing tracks and satisfying the minimum bend radius.
+Staggered Port Access Region Reservation.
+To prevent waveguides from blocking ports, a region in front of each port is reserved along its orientation (Fig.
+4(a)).
+For multiport devices such as MMIs, assigning the same reserved length to all ports can still cause blockage, while closely spaced parallel waveguides hinder crossing insertion (Fig.
+4(c)), since a crossing requires a minimum footprint and cannot be placed consecutively without sufficient spacing.
+To address this, we introduce the staggered port access region strategy, ensuring adjacent access points are spaced larger than the crossing footprint, leaving sufficient room for accessibility.
+The reserved length depends on the port’s order among ports of the same orientation, as well as crossing size and bend radius.
+Ports are grouped by orientation and sorted by coordinates along the perpendicular axis, each assigned an index .
+The reserved length for port is calculated by:
+where is the number of ports in the group.
+This yields a staggered, mountain-shaped profile that prevents inner ports from being blocked and facilitates escape.
+If routing fails, the reserved length is adjusted according to the updated routing order in Section IV-B.
 III-B Port-Group-based Net Order
@@
 Details are provided in Section IV.
-III-C Non-Manhattan Waveguide Routing with Curvy-Aware A∗ Search
+III-C Non-Manhattan Waveguide Routing with Curvy-Aware A∗
 Unlike Manhattan-style VLSI routing, PICs require non-Manhattan paths to support smooth curves that reduce bending and insertion loss.
@@
 For example, for an MS node oriented at 0∘, its directly adjacent neighbor is 1 grid unit away, while the steps to reach the 90∘ and 45∘ neighbors are calculated as:
-The ceiling function guarantees that there is enough spacing to accommodate bending structures.
+Here, and represent the horizontal and vertical grid distances required to reach a neighbor at a 90∘ turn.
+The ceiling function guarantees that the bend radius constraint is met.
 Note that unlike 45∘ diagonal neighbors in traditional 8-way A∗ search, where is equal to , our neighbors are the endpoint of the 45-degree curves, where and are not equal.
@@
 The total cost of a node in A∗ search is given by the standard formulation: , where denotes the accumulated cost from the source node to the current node , and is the heuristic estimate of the cost from node to the target node .
-The cost consists of two components: the insertion loss along the path to the current node, denoted as and calculated based on Eq.
-(1), and a group-based congestion penalty (GCP), denoted as to save routing resources for a group of nets:
+The cost includes an insertion-loss-inspired term , adapted from Eq.
+(1) with a tunable weight and a group-based congestion penalty (GCP), denoted as to save routing resources for a group of nets.
+By adjusting the weight parameters, users can flexibly balance between waveguide length and the number of crossings to meet specific design objectives:
 where is a penalty coefficient that discourages routing too close to blockages or previously routed waveguides, preserving routing resources for unrouted nets within the same routing group, and is the number of grids that occupied by others in the check region as shown in Fig.
@@
 As more nets are routed, decreases to avoid consuming extra space.
-Since our method allows non-Manhattan routing, the shortest path is often diagonal.
-To ensure the A∗ heuristic remains valid and does not overestimate the cost, we define a customized heuristic as shown in Eq.
-(5):
+Since our method allows non-Manhattan routing, the shortest path is often diagonal. We define a customized heuristic as shown in Eq. (6).
 where and represent the minimum and maximum differences in the x and y coordinates between the current node and the target node , respectively.
-Additionally, a penalty term proportional to the 45∘ bending loss, , is added when both and are non-zero, as this scenario requires at least one 45∘ bend.
-This penalty encourages paths that end in orientations better suited for port connection.
+When , we have , and reduces to the standard Manhattan case.
+When , it indicates that the subsequent path search must contain at least one bend with an angle no smaller than .
+Therefore, we incorporate the loss of a bend into , which does not overestimate the cost and thus the A∗ router remains admissible, and penalizes the bend and encourages paths that end in orientations suited for port connection.
 III-C5 Waveguide Instantiation
@@
 IV-A Hierarchical Netlist Tree Construction
-PIC netlists often exhibit inherent hierarchical structure, where complex systems are composed of repeated functional modules as shown in Fig.
+PIC netlists often exhibit inherent hierarchical structure, where complex systems are composed of repeated functional modules, as shown in Fig.
 11.
 Leveraging this structure is crucial for scalable physical design.
-To exploit such hierarchy, we first construct a Hierarchical Tree based on the invocation relationships between modules and the overall circuit.
+To exploit such hierarchy, we construct a Hierarchical Tree by parsing the nested module instantiations defined in the YAML-format netlist, which is similar in style to a SPICE netlist and inherently preserves hierarchical relationships.
+In this tree, the root node corresponds to the top-level module of the circuit, the leaf nodes represent primitive photonic components, and the internal nodes represent intermediate-level submodules.
+The resulting tree has arbitrary fanout and depth, depending on the circuit topology.
+The construction has linear time complexity of , where is the total number of unique modules and components.
 This hierarchical representation enables global routing planning at multiple levels.
-We then perform routing in a bottom-up manner, completing routing for each level of modules progressively.
+We then perform bottom-up routing for each level of modules progressively.
+Within each level, the priority of each routing group is determined according to the group priority score defined in Subsection III-B.
 This multi-level routing approach helps avoid conflicts between different modules, as routing is localized within each submodule before higher-level integration.
@@
 Offset Neighbor.
-Although the algorithm naturally discourages excessive bending through its bend cost heuristic, it may still introduce unnecessary bends when ports are slightly misaligned.
-Especially in tight spaces, such misalignment can trigger multiple bends, increasing insertion loss and risking routing failure, as shown in Fig.
-12(a).
-To address this issue, we incorporate additional offset neighbors to avoid such detours under specific conditions: when the current node is close to the target node (within four times the bending radius), and the orientation differs by 180 degrees from the target node.
-These offset neighbors are designed to correct small misalignments (less than the bending radius) between the source and target ports.
-The offset neighbor locations are computed analytically based on the predefined bending radius and grid size, ensuring that the resulting path remains smooth and physically realizable using sine-bend waveguides.
-From an algorithmic perspective, the introduction of offset neighbors simplifies port access, reduces redundant neighbor exploration caused by bending detours, and thereby improves the search efficiency.
+Although the algorithm discourages excessive bends via its bend-cost heuristic, it may still add unnecessary bends when ports are slightly misaligned.
+In tight spaces, such misalignment can cause multiple bends, raising insertion loss and risking routing failure (Fig.
+12(a)).
+To address this, we add offset neighbors under specific conditions: when the current node is within four bending radii of the target and their orientations differ by 180 degrees.
+These neighbors correct small misalignments (below the bend radius) between source and target ports.
+Their locations are computed analytically from the bend radius and grid size, ensuring smooth, fabricable sine-bend waveguides.
+From an algorithmic perspective, offset neighbors simplify port access, cut redundant detours, and improve search efficiency.
 Preserved Crossing Space.
@@
 The penalty is defined as:
-where is a user-defined penalty coefficient and denotes the minimum width required to insert a valid crossing structure.
+where is a user-defined penalty and is the minimum width required to insert a valid crossing.
 The penalty is applied over a -wide region around the net segment, discouraging other nets from occupying this reserved space.
 IV-C Reduce Routing Conflict via Routing Order Optimization
-Although the port region allows access points to be staggered, reducing conflicts during access, an inappropriate access order may easily result in the target ports being blocked by routed waveguides and other port regions.
-Especially for multi-port components, the intra-group routing order is critically important.
+Although port regions allow staggered access points to reduce conflicts, an improper order may still block target ports with routed waveguides or other regions.
+For multi-port components, the intra-group routing order is thus crucial.
 As shown in Fig.
-13(a), both the orientation and spatial location of ports significantly influence the routing order, as they jointly determine the accessibility and potential for routing conflicts among nets.
-To address the intra-group routing order, we determine the routing order by the spatial relationship between source ports and target ports within the group, as shown in Fig.
-13(b).
-Let the set of source ports be denoted as .
-We compute their geometric mean coordinates.
-If each port has a position , the average position is:
-After obtaining the mean location of the source ports, we determine the routing order based on the Manhattan distance from each target port to .
-Target ports that are closer to the average source location are assigned higher routing priority.
-This approach can effectively resolve most routing conflict issues, particularly in multiport components.
+13(a), both orientation and spatial location of ports strongly influence order, as they determine accessibility and potential conflicts.
+To address this, we derive the order from the spatial relation between source and target ports (Fig. 13(b)). Let the source port set be with positions . Their mean coordinates are:
+Routing order is then based on Manhattan distance from each target port to , with closer ports given higher priority.
+This method effectively resolves most conflicts, especially in multiport components.
 Routing Order Refinement.
 It is difficult to guarantee an optimal routing order.
-In particular, when nets are routed independently without group information, the aforementioned method becomes inapplicable.
-Hence, it is necessary to dynamically adjust the order during the routing process.
-An inappropriate order may lead to ports being blocked by previously routed waveguides, resulting in escape or access failures.
-As illustrated in Fig.
-14, since net1, net2, and net3 are routed independently, net2 fails because it cannot access the port due to net1’s port region and net3’s waveguide blocking its path.
-To address routing failures, we enhance the rip-up and re-route scheme by not only removing nets that directly conflict with the failed path but also ripping up those located in front of the failed net’s ports to alleviate congestion.
+When nets are routed independently without group information, the method above is inapplicable, making dynamic order adjustment necessary.
+An improper order may block ports with earlier waveguides, causing escape or access failures.
 As shown in Fig.
-14, we scan along the port orientation and identify the conflict order and names of the conflicting nets (starting from ), which serve as guidance to adjust the routing order.
-Higher-priority nets are routed first to improve escape success, but may block others’ access.
-To avoid this, we dynamically extend the target port regions of high-priority nets, shifting their access points and reducing the risk of subsequent blocking.
+14, routing net1, net2, and net3 independently leads to net2’s failure, as net1’s port region and net3’s waveguide block its path.
+To address routing failures, we enhance rip-up and re-route by removing not only nets directly conflicting with the failed path but also those in front of the failed net’s ports to ease congestion.
+As shown in Fig.
+14, we scan along the port orientation, identify conflict order and conflicting nets (starting from ), and use them to adjust routing order.
+For instance, if net3 is blocked by net2, and net2 by net1, we merge them and uniformly update the routing order from 3 to 1.
+When net3 and net2 are both blocked by net1 and their order is indistinguishable, we sort them by net names.
+Although higher-priority nets are routed first, they may still block others.
+To prevent this, we dynamically extend high-priority target ports.
+Specifically, the port length is extended by , where is the number of merged nets, the order, and the bend radius, shifting access points and reducing blocking risk.
 Our extended conflict-resilient hierarchical routing scheme can effectively enhance routability even under tight layout constraints and help accelerate the routing process.
@@
 | GWOR_3232_C | 33 | 64 | – | 6000 6000 |
-| Bennes_1616_C | 224 | 416 | 88 | 2200 1300 |
-| Bennes_3232_C | 576 | 1024 | 416 | 3500 2500 |
+| Benes_1616_C | 224 | 416 | 88 | 2200 1300 |
+| Benes_3232_C | 576 | 1024 | 416 | 3500 2500 |
 PIC Intermediate Representation.
@@
 This benchmark emphasizes routing conflict resolution via pure spatial reasoning.
-Bennes [45]: Based on Mach-Zehnder interferometers, it adopts a structured, hierarchical Benes topology.
+Benes [38]: Based on Mach-Zehnder interferometers, it adopts a structured, hierarchical Benes topology.
 It exhibits extremely high crossing density, where a single net may require tens of crossings, challenging both crossing reservation and congestion-aware routing strategies.
@@
 Table II (ISPD version) includes spacious, moderately complex layouts, while Table III features more compact benchmarks with higher component density, smaller die sizes, and more crossings, posing greater routing challenges and better reflecting real-world photonic layout constraints.
-The bending radius is set to 60 for Light and GWOR and 5 for others.
+For the Light and GWOR benchmarks, the bending radius, waveguide width, and grid resolution are set to 60 m, 2 m, and 50 m, respectively, to emulate the larger bend radii of the SiN platform.
+Since these circuits provide abundant routing resources, a larger grid size is adopted to accelerate the search process.
+For the other benchmarks, these values are set to 5 m, 0.5 m, and 2 m, respectively, reflecting the tighter bends of the SiPh platform.
+By default, the crossing size is fixed at 10 m 10m.
 | Benchmarks | PROTON [12] (Adaptive crossing penalty) | LiDAR 1.0 [39] | LiDAR 2.0 | ||||||||||||
@@
 The proposed photonic detailed routing framework is implemented in Python based on GDSFactory 8 [31] libraries.
-All evaluations are conducted on a on a Linux server with a 128-core AMD EPYC 7763 CPU and 512-GB memory.
+All evaluations are conducted on a Linux server with a 128-core AMD EPYC 7763 CPU and 512-GB memory.
 Note that runtime and results may slightly differ from the ISPD version due to platform and GDSFactory version variations.
@@
 Device insertion losses used for critical path IL evaluation are listed in Table V, with bending loss proportional to angle.
+A∗ Search Parameters.
+Referring to the parameters in Table V and scaling them to a unified micrometer unit, the base propagation, bending, and crossing weights are set to 1.5, 50, and 5000, respectively.
+On top of these, we further adjust the weights according to the characteristics of each benchmark.
+For the photonic tensor core, where crossings are inherent and unavoidable, we set the crossing weight equal to the bending weight to speed up search.
+For the optical switch, the crossing weight remains 5000, while the bending weight is also set to 5000 to reduce bends; in long waveguides with large radii, shorter wirelength can offset added bend loss.
+The group congestion penalty (GCP) parameter is determined by the detour distance, as saving routing resources through GCP may conflict with wirelength minimization.
+We set this value to 500, corresponding to reserving approximately 200 m of spacing.
+The crossing spacing coefficient is set similarly, but its checking range is limited to a single crossing.
 Baselines.
@@
 Despite the Clements circuit having a regular topology with no inherent crossings and relatively spacious layouts, insufficient planning and lack of reserved access space lead to unnecessary crossings and DRVs in PROTON.
-These routing conflicts become more severe in ADEPT benchmarks, which features dense multi-port MMIs and a large number of crossings.
+These routing conflicts become more severe in ADEPT benchmarks, which feature dense multi-port MMIs and a large number of crossings.
 By incorporating Accessibility-Enhanced Port Assignment and Port-Group-Based Net Ordering, LiDAR 2.0 effectively mitigates these conflicts and reduces unnecessary rip-up and reroute.
@@
 Light benchmark provides abundant routing resources and is primarily used to evaluate the router’s ability to optimize insertion loss and crossings.
-Notably, there exists a counter-intuitive trade-off between #CR and WL, minimizing #CR may lead to long detours that increase propagation loss, resulting in a higher overall .
+Notably, there exists a counter-intuitive trade-off between #CR and WL; minimizing #CR may lead to long detours that increase propagation loss, resulting in a higher overall .
 Except for the Light_b case, LiDAR achieves the lowest across all other cases while maintaining crossing-optimal (#CR=0), DRV-free layouts.
@@
 | TeMPO_88_C | 12 | 1.73 | 18.48 | 14 | 95 | 10 | 1.90 | 17.33 | 5 | 147 | 7 | 2.83 | 15.86 | 0 | 44 |
-| TeMPO_1616_C | 27 | 3.47 | 34.58 | 35 | 251 | 23 | 3.86 | 32.27 | 24 | 1571 | 15 | 5.48 | 28.31 | 0 | 119 |
+| TeMPO_1616_C | 27 | 3.47 | 34.58 | 35 | 251 | 23 | 3.86 | 32.27 | 24 | 1571 | 15 | 5.48 | 28.31 | 0 | 169 |
 | TeMPO_3232_C | 49 | 5.91 | 62.57 | 113 | 1125 | 46 | 6.11 | 60.45 | 57 | 8057 | 31 | 10.96 | 53.27 | 0 | 425 |
@@
 | GWOR_3232_C | 62 | 7.22 | 35.26 | 126 | 60 | 56 | 6.81 | 31.26 | 44 | 45 | 44 | 7.82 | 25.40 | 0 | 11 |
-| Bennes_1616_C | 33 | 4.42 | 22.69 | 76 | 721 | 26 | 4.28 | 18.84 | 3 | 175 | 23 | 3.57 | 17.15 | 0 | 33 |
-| Bennes_3232_C | 56 | 9.24 | 36.76 | 365 | 5117 | 68 | 9.08 | 42.39 | 78 | 1690 | 58 | 6.51 | 36.56 | 0 | 251 |
+| Benes_1616_C | 33 | 4.42 | 22.69 | 76 | 721 | 26 | 4.28 | 18.84 | 3 | 175 | 23 | 3.57 | 17.15 | 0 | 33 |
+| Benes_3232_C | 56 | 9.24 | 36.76 | 365 | 5117 | 68 | 9.08 | 42.39 | 78 | 1690 | 58 | 6.51 | 36.56 | 0 | 251 |
 | Geo-mean | - | 5.39 | 29.21 | - | 800 | - | 5.27 | 27.05 | - | 1078 | - | 5.61 | 24.58 | - | 155 |
 | Ratio | - | 1 | 1 | - | 1 | - | 0.97 | 0.92 | - | 1.33 | - | 1.04 | 0.84 | - | 0.19 |
-We further evaluate the routing performance on compact benchmarks characterized by higher component density and more topological crossings.
-For these cases, only a single round of rip-up and reroute is applied, as limited layout space makes additional iterations ineffective.
-Figure 15 shows the final routed layouts generated by LiDAR 2.0.
-As shown in Table VI, LiDAR 2.0 achieves nearly DRV-free results across all benchmarks, with an average 16% reduction in insertion loss and a 5.16 speedup over PROTON, and a 6.95 speedup and 9% lower insertion loss compared to LiDAR 1.0.
-On the TeMPO benchmark, PROTON shows a much lower runtime due to complete port blocking and early termination without successful routing.
-On the compact cases, LiDAR 1.0 also shows increased DRVs and runtime, especially on newly added cases.
-In contrast, LiDAR 2.0 adopts a hierarchical routing strategy that enables subcircuit reuse in designs such as TeMPO and Bennes, significantly improving runtime.
-Building on this, intra-group routing order optimization and proactive crossing space preservation further reduce conflicts in multiport components, improve solution quality, and avoid unnecessary search.
-As a result, LiDAR 2.0 successfully achieves DRV-free routing even on high-crossing-density Bennes benchmarks.
-For GWOR, which lacks group information, LiDAR 2.0 still generates DRV-free results by dynamically refining the routing order and port regions during rip-up and reroute.
-Figure 16 presents the runtime breakdown across various routing stages on Bennes_3232.
-The substantial speedup achieved by LiDAR 2.0 stems from its ability to minimize routing conflicts, eliminate redundant searches and RR iterations, and hierarchically reuse routing results.
+We further evaluate routing on compact benchmarks with higher component density and more crossings.
+Only one rip-up and reroute round is applied, as limited space makes further iterations ineffective.
+Figure 15 shows the final layouts by LiDAR 2.0.
+As in Table VI, LiDAR 2.0 delivers nearly DRV-free results with an average 16% IL reduction, 5.16 speedup over PROTON, and 6.95 speedup with 9% lower IL vs.
+LiDAR 1.0.
+On TeMPO, PROTON runs faster only because complete port blocking forces early termination.
+For compact cases, LiDAR 1.0 shows more DRVs and runtime, especially on new benchmarks.
+In contrast, LiDAR 2.0’s hierarchical routing reuses subcircuits (TeMPO, Benes), greatly improving runtime.
+Intra-group ordering and proactive crossing-space reservation further reduce conflicts, enhance quality, and avoid redundant search.
+Thus, LiDAR 2.0 achieves DRV-free routing even on high-crossing Benes.
+For GWOR, lacking group data, LiDAR 2.0 still succeeds by dynamically refining order and port regions during reroute.
+Figure 16 shows the runtime breakdown on Benes_3232. The speedup of LiDAR 2.0 comes from reducing conflicts, eliminating redundant search and RR iterations, and hierarchically reusing results.
 VI-C Results on Different Crossing Sizes and Bend Radii
@@
 | TeMPO_1616_L | 15 | 4.86 | 28.22 | 0 | 132 | 15 | 4.79 | 28.21 | 0 | 136 | 15 | 4.71 | 28.19 | 0 | 139 |
-| Bennes_1616_C | 23 | 3.57 | 17.15 | 0 | 33 | 23 | 3.45 | 17.13 | 0 | 34 | 23 | 3.34 | 17.15 | 0 | 34 |
-| Bennes_1616_L | 22 | 3.70 | 16.68 | 0 | 34 | 22 | 3.57 | 16.62 | 0 | 28 | 23 | 3.49 | 17.23 | 0 | 35 |
-| Metrics | GWOR_3232 | Clements_1616_C | ADEPT_1616_C | Bennes_1616_C | TeMPO_1616_C | ||||||||
+| Benes_1616_C | 23 | 3.57 | 17.15 | 0 | 33 | 23 | 3.45 | 17.13 | 0 | 34 | 23 | 3.34 | 17.15 | 0 | 34 |
+| Benes_1616_L | 22 | 3.70 | 16.68 | 0 | 34 | 22 | 3.57 | 16.62 | 0 | 28 | 23 | 3.49 | 17.23 | 0 | 35 |
+| Metrics | GWOR_3232 | Clements_1616_C | ADEPT_1616_C | Benes_1616_C | TeMPO_1616_C | ||||||||
 | =10 | =20 | =40 | =80 | =100 | =10 | =15 | =10 | =15 | =5 | =10 | =5 | =10 | |
-| WL | 7.77 | 7.77 | 7.79 | 7.85 | 8.06 | 3.49 | 3.33 | 8.55 | 8.52 | 3.57 | 3.76 | 2.83 | 3.07 |
-| #CR | 44 | 44 | 44 | 44 | 44 | 0 | 3 | 21 | 22 | 23 | 23 | 7 | 15 |
-| 26.40 | 25.40 | 25.40 | 25.41 | 25.44 | 26.54 | 26.25 | 21.06 | 21.62 | 17.15 | 17.32 | 15.86 | 27.90 | |
+| WL | 7.77 | 7.77 | 7.79 | 7.85 | 8.06 | 3.49 | 3.33 | 8.55 | 8.52 | 3.57 | 3.76 | 5.48 | 5.41 |
+| #CR | 44 | 44 | 44 | 44 | 44 | 0 | 3 | 21 | 22 | 23 | 23 | 15 | 15 |
+| 26.40 | 25.40 | 25.40 | 25.41 | 25.44 | 26.54 | 26.25 | 21.06 | 21.62 | 17.15 | 17.32 | 28.32 | 27.90 | |
 | #DRV | 0 | 0 | 0 | 0 | 3 | 0 | 3 | 0 | 2 | 0 | 0 | 0 | 0 |
-| Time (s) | 10 | 10 | 10 | 14 | 16 | 96 | 111 | 298 | 339 | 33 | 33 | 44 | 172 |
+| Time (s) | 10 | 10 | 10 | 14 | 16 | 96 | 111 | 298 | 339 | 33 | 33 | 169 | 172 |
 Different PIC technologies impose varying constraints on crossing sizes and bending radii.
 Table VII and Table VIII show the routing results of LiDAR 2.0 under different crossing sizes and bend radii.
+We synthetically stretch the device only to increase the routing difficulty and test our router’s capability, regardless of its actual IL value.
 Benchmarks with an L suffix indicate larger component pitch.
@@
 For silicon photonics, where the bending radius is typically 5–10 , LiDAR 2.0 performs reliably.
-Feasible solutions remain available at 15 despite minor DRVs, except for the TeMPO and Bennes benchmarks, where no valid routing solution exists due to insufficient spacing.
+Feasible solutions remain available at 15 despite minor DRVs, except for the TeMPO and Benes benchmarks, where no valid routing solution exists due to insufficient spacing.
 VI-D Discussion
+| Metrics | Light_a | ADEPT_44 | ||
+| Manual | LiDAR 2.0 | Manual | LiDAR 2.0 | |
+| #CR | 0 | 0 | 10 | 10 |
+| WL (mm) | 31.89 | 29.99 | 2.30 | 2.33 |
+| 7.85 | 7.61 | 12.95 | 13.14 | |
+| DRV | 0 | 0 | 0 | 0 |
+| Time | 2h | 103s | 2h | 26s |
 | Metrics | High Crossing Cost | Low Crossing Cost | ||
-| w/o GCP | LiDAR | w/o GCP | LiDAR | |
+| w/o GCP | LiDAR 2.0 | w/o GCP | LiDAR 2.0 | |
 | #CR | 6 | 0 | 5 | 5 |
@@
 | Time (s) | 129 | 103 | 261 | 197 |
+| Benchmarks | w/o Staggered offset | w/o crossing space | w/o group-based net order | |||
+| #DRV | #DRV | #DRV | ||||
+| ADEPT_1616_C | 18.95 | 37 | 19.97 | 6 | 21.63 | 4 |
+| ADEPT_1616_L | 19.13 | 33 | 20.34 | 3 | 21.95 | 3 |
+| TeMPO_1616_C | 28.49 | 27 | 28.31 | 0 | 31.87 | 16 |
+| TeMPO_1616_L | 28.72 | 26 | 28.22 | 0 | 32.64 | 16 |
+| Benes_1616_C | 14.62 | 21 | 16.24 | 32 | 16.83 | 2 |
+| Benes_1616_L | 14.37 | 21 | 16.93 | 30 | 17.21 | 4 |
 | Benchmarks | w/o Offset Neighbor | w/ Offset Neighbor | ||||
@@
 | TeMPO_3232_C | 53.94 | 0 | 673 | 53.27 | 0 | 425 |
+Manual Routing Result.
+We conducted a manual routing comparison on small circuits, Light_a and ADEPT_44, as shown in TABLE IX.
+It can be observed that the routing quality of LiDAR 2.0 is comparable to that of manual routing, and even achieves lower insertion loss on Light_a.
+However, manual routing required approximately two hours to complete for these small circuits, and as the circuit scale increases, manual efforts become impractical.
+Spacing Reservation.
+We propose several techniques for reserving spacing, including staggered offset port region, reserved crossing spacing, group congestion penalty, and group-based net ordering to mitigate routing conflicts in multiport components.
+TABLE XI highlights their impact on representative circuits.
+It can be observed that the staggered offset port region is most critical: removing it causes many routing failures.
+Without the crossing space penalty, the Benes circuit suffers heavily due to consecutive crossings, while TEMPO is less affected since its crossings are independent.
+Disabling group-based net ordering, however, triggers numerous DRVs in TEMPO because of high-fanout components.
+We also assess the proposed GCP in guiding crossing optimization under different crossing weights (Table X).
+With larger , the router avoids crossings to reduce maximum insertion loss (); with smaller , it favors shorter paths even with extra crossings, further improving .
+Without GCP, more crossings arise as increases, since earlier nets congest low-crossing paths.
+With GCP, routing conflicts drop notably, allowing efficient paths with fewer crossings and lower IL.
+In Fig.
+17, flexibly balances waveguide length and crossings to meet performance needs.
 Crossing-Waveguide Optimization.
 As shown in Table IV (PROTON [12] vs.
-LiDAR 2.0), our proposed crossing optimization strategy introduces an extra runtime penalty on the Light benchmark, but it leads to higher solution quality as it reduces the crossings between IOs and the switch given the high crossing penalty.
-Port-Group-based Congestion Penalty (GCP).
-We evaluate the effectiveness of our proposed group-based congestion penalty in guiding crossing optimization under different crossing insertion loss settings, as shown in Table IX.
-Specifically, we compare scenarios with high crossing IL () and low crossing IL ().
-When the crossing IL is high, our method effectively avoids crossings to minimize the maximum insertion loss ().
-In contrast, under low crossing IL, it prefers shorter waveguide paths with more crossings to further optimize .
-Without the group-based penalty, however, the router tends to use more crossings as increases.
-This is because, under congestion from previously routed nets, it struggles to find low-crossing paths.
-By introducing the group-based penalty, routing conflicts are significantly reduced, enabling more efficient decisions with fewer crossings and lower overall IL.
-As illustrated in Fig.
-17, the coefficient functions as a flexible control parameter, allowing users to balance between waveguide length and the number of crossings based on specific PIC performance requirements.
+LiDAR 2.0), our proposed crossing optimization strategy introduces an extra runtime penalty on the Light benchmark, but it leads to higher solution quality as it reduces the crossings between IOs and the switch, given the high crossing penalty.
 Offset Neighbor.
-Offset neighbors help eliminate unnecessary bending detours caused by port misalignment in PICs, thereby reducing both routing conflicts and search overhead.
-To further demonstrate their effectiveness, we evaluate the router on a compact TeMPO_C+ variant with x/y pitch reduced from 25/50 to 20/15 , as shown in Table X.
-In such highly constrained layouts, detours without offset neighbors often result in design rule violations.
-Even in more spacious cases, enabling offset neighbors consistently leads to faster runtime compared to routing without them.
+Offset neighbors help eliminate unnecessary bending detours caused by port misalignment, thereby reducing routing conflicts and search overhead.
+We evaluate them on a compact TeMPO_C+ variant with x/y pitch reduced from 25/50 to 20/15 (Table XII).
+In such constrained layouts, routing without offset neighbors often causes rule violations.
+Even in spacious cases, enabling them yields consistently faster runtime.
 VII Conclusion
-In this work, we first introduce LiDAR, an open-source automated detailed routing tool specifically designed for photonic integrated circuits.
-LiDAR features a non-Manhattan curvy-aware A∗ search engine with accessibility-enhanced port assignment, adaptive crossing insertion, congestion-aware group-based net ordering and objective, and crossing optimization scheme to handle unique PIC routing constraints while optimizing critical path insertion loss.
-To further alleviate routing conflicts and improve efficiency, we extend LiDAR into a more scalable and conflict-resilient framework, LiDAR 2.0, by introducing hierarchical routing reuse, intra-group net ordering, offset-neighbor search, and crossing-space preservation.
-These enhancements substantially improve routability under tight layout constraints and high crossing density.
-LiDAR 2.0 achieves nearly DRV-free layouts across challenging benchmarks, delivering up to 9% lower insertion loss and 6.95 speedup compared to LiDAR 1.0, advancing the automation frontier for large-scale PIC physical design.
+In this work, we present LiDAR, an open-source detailed router for PICs.
+It features a non-Manhattan curvy-aware A∗ engine with enhanced port assignment, adaptive crossing insertion, congestion-aware group-based net ordering, and crossing optimization to address unique PIC constraints while minimizing critical path insertion loss.
+We further extend it into a scalable, conflict-resilient framework, LiDAR 2.0, by adding hierarchical routing reuse, intra-group ordering, offset-neighbor search, and crossing-space preservation.
+These enhancements markedly improve routability under tight layouts and high crossing density.
+LiDAR 2.0 achieves nearly DRV-free layouts on challenging benchmarks, with up to 9% lower IL and 6.95 speedup over LiDAR 1.0, advancing automation for large-scale PIC design.